# Large Language Models for Cyber Security: A Systematic Literature Review

### v4

@@ 3.5. Application of LLMs in Blockchain Security
RQ1 - Summary (1) We have divided cybersecurity tasks into six domains: software and system security, network security, information and content security, hardware security, and blockchain security. We have summarized the specific applications of LLMs in these domains. (2)We discussed 29 cybersecurity tasks and found that LLMs are most widely applied in the field of software and system security, with 119 papers covering 11 tasks.
@@ 4.2. Trend Analysis
RQ2 - Summary (1) We have gathered papers utilizing over 30 distinct LLMs for cybersecurity tasks. These LLMs have been categorized into three groups based on their underlying architecture or principles: encoder-only, encoder-decoder, and decoder-only LLMs. (2) We analyzed the trend in employing LLMs for cybersecurity tasks, revealing that decoder-only architectures are the most prevalent. Specifically, over 15 LLMs fall into the decoder-only category, and a total of 133 papers have investigated the utilization of decoder-only LLMs in cybersecurity tasks.
@@ 5.1. Fine-tuning LLMs for Security Tasks
Beyond standard supervised fine-tuning, Reinforcement Learning (RL) based strategies are gaining attention for aligning LLMs more closely with specific security objectives or leveraging their outputs to guide automated processes (Ouyang et al., 2022; DeepSeek-AI et al., 2025; François-Lavet et al., 2018). These approaches often involve optimizing an agent’s policy using reward signals derived from task execution or human feedback, sometimes incorporating LLM-generated insights into the reward function (Tang et al., 2024). For instance, in web application security, RL can fine-tune an LLM to generate more effective web fuzzing inputs by rewarding payload sequences that successfully bypass defenses or trigger desired responses (Liang et al., 2024). Similarly, RL techniques can enhance log analysis by optimizing models based on the accuracy or utility of the identified anomalies (Han et al., 2023). In the domain of fuzzing complex systems like JavaScript engines, CovRL utilizes coverage-guided RL: it processes execution coverage using methods like TF-IDF to assign higher importance to rarely executed paths, then uses this weighted coverage to generate a reward signal, thereby fine-tuning an LLM mutator to produce mutations more likely to explore these less-covered, potentially bug-rich areas (Eom et al., 2024). For automated program repair, RePair leverages RL with process-based feedback; instead of just learning from fixed code examples, the LLM’s policy for generating patches is directly rewarded based on the outcome of executing the generated code against test suites, thus encouraging the model to learn repair strategies that lead to functional correctness (Zhao et al., 2024). In a different application connecting LLMs and RL, CrashTranslator utilizes LLMs first to analyze stack traces and compute semantic similarities with UI elements; this information then shapes the reward function for an RL agent that learns an optimized UI exploration policy to effectively reproduce mobile application crashes (Huang et al., 2024).
@@ 5.3. External Augmentation
External tools Analysis results from specialized tools serving as auxiliary inputs or validate the result with external tools. Static analysis tools, penetration testing tools, Rule-Based Reasoning Engine (Happe et al., 2023) (Alrashedy and Aljasser, 2024) (Armengol-Estapé et al., 2024) (Cao et al., 2024)(Nazzal et al., 2024) (Xu et al., 2025)(Chen and Jabbarvand, 2024) (Zhang et al., 2024c)(Zhang et al., 2024d) (Huang et al., 2024)(Zhang and Zhang, 2024) (Ahmad et al., 2024a)
@@
Inter-model interaction Introducing multiple models (which can be LLMs or other models) to collaborate and interact. Multiple LLMs collaboration, graph neural networks, gGAN (Cai et al., 2024a) (Yang et al., 2024a)(Tang et al., 2023) (Nazzal et al., 2024)(Huang et al., 2024) (Mündler et al., 2024)(Wang et al., 2024b)
Rebroadcasting Applicable to multi-step tasks, broad-casting the output results of each step iteratively as part of the prompt for the next step. Difficulty-based patch example replay, variables’ name propagation (Xu et al., 2023) (Yuan et al., 2022)(Xu et al., 2025)
@@
We summarized the external augmentation techniques combined with LLMs in previous studies, as shown in Table 7, identifying 8 distinct categories. The first augmentation technique is feature augmentation. The effectiveness of LLMs in handling downstream tasks heavily relies on the features included in the prompts. We have observed that many studies employing LLMs for security tasks extract contextual relationships or other implicit features from raw data and integrate them with the original data to customize prompts. These implicit features encompass descriptions of vulnerabilities (Zhang et al., 2023f), bug locations (Jin et al., 2023), threat flow graphs (Liu et al., 2023c), and more. Incorporating these implicit features alongside raw data leads to enhanced performance compared to constructing prompts solely from raw data. The next augmentation technique is external retrieval. External knowledge repositories can mitigate the hallucinations or errors arising from the lack of domain expertise in LLMs. LLMs can continually interact with external knowledge repositories during pipeline processing and retrieve knowledge relevant to security tasks to provide superior solutions (Perrina et al., 2023). Next are external tools; rule-based external tools can serve as specialized external knowledge repositories, and in addressing security tasks, LLMs can utilize results from external tools (like static analyzers or formal methods) to rectify their outputs, thereby avoiding redundancy and errors(Happe et al., 2023; Alrashedy and Aljasser, 2024). The fourth augmentation technique is task-adaptive training. Existing studies adopt various training strategies (beyond standard fine-tuning discussed in Section 5.1), like contrastive learning or distillation, to strengthen LLMs’ adaptability to complex security tasks, enabling them to generate more targeted outputs (Chen et al., 2020; Tang et al., 2023; Fang et al., 2023). The fifth augmentation technique, inter-model interaction, has garnered significant attention when a single LLM may struggle to handle complex and intricate tasks. Decomposing the pipeline process and introducing multiple LLMs for enhanced performance have been explored (Yang et al., 2024a). This approach leverages collaboration and interaction among models to harness the underlying knowledge base advantages of each LLM. When a single interaction is insufficient to support LLMs in tasks such as variable name recovery or generating complex program patches  (Xu et al., 2023; Yuan et al., 2022), rebroadcasting is used: constructing prompts for LLMs multiple times continuously where the output results of each step are broadcast iteratively as part of the prompt for the next step. This helps reinforce the contextual relationship between each interaction, thereby reducing error rates. The next augmentation technique is post-processing, where LLMs’ outputs are validated or processed for certain security tasks requiring specific types of output (Tihanyi et al., 2023). This process helps mitigate issues such as hallucinations arising from the lack of domain knowledge in LLMs (Chen et al., 2023c). The last strategy is workflow integration, which involves integrating LLMs as components within specific task workflows, incorporating various forms of interactive feedback. This feedback, such as coverage metrics guiding fuzzing strategies (Wang et al., 2024f; Lyu et al., 2024), self-check reasoning steps (Wang et al., 2024f), or outputs being integrated into human rating systems (Thomas et al., 2024), allows the workflow’s context and objectives to steer the LLM’s behavior.
@@
RQ3 - Summary (1) We summarize the domain-specific techniques used in previous research to apply LLMs to security tasks, including prompt engineering, fine-tuning, and external augmentation. (2) Prompt engineering is the most widely used domain technique, with almost all 185 papers employing this approach. Fine-tuning techniques were used in 31% of the papers, while task-specific external augmentation techniques were adopted in 36% of the papers. (3) We categorize and discuss the fine-tuning and external augmentation techniques mentioned in these papers, and analyze their relevance to specific security tasks.
@@ 6.2. Types of Datasets
Buggy code and bug reports 4 (Zhang et al., 2024c) (Huang et al., 2024) (Wang et al., 2024c) (Tang et al., 2024)
@@
Decompiled code and variable names 1 (Xu et al., 2025)
@@
Combined datasets, integrating multiple data types, represent a growing approach (23 instances identified). This strategy often pairs code artifacts with relevant natural language text, such as combining vulnerable code with vulnerability descriptions (Liu et al., 2023d; Chen et al., 2023c), buggy code with bug reports (Zhang et al., 2024c; Huang et al., 2024), or source code with program documentation (Zhang et al., 2024e). The key advantage lies in leveraging the LLM’s ability to process both structured code and unstructured text simultaneously (Wang et al., 2024a). By providing natural language context (e.g., descriptions of what a vulnerability entails or what a function should do), these combined datasets help LLMs achieve a deeper semantic understanding of the code during fine-tuning or prompting, leading to more accurate vulnerability detection, better code repair suggestions, or more effective code analysis compared to using code data alone.
@@ 6.4. Data Augmentation using LLMs
RQ4 - Summary (1) Based on different data sources, datasets are categorized into four types: open-source datasets, collected datasets, constructed datasets, and industrial datasets. The use of open-source datasets is the most common, accounting for approximately 42% in the 173 papers explicitly mentioning dataset sources. Collected datasets and constructed datasets are also popular, reflecting the lack of practical data in LLM4Security research. (2) We categorize all datasets into three types: code-based, text-based, and combined. Code-based and text-based types are the most commonly utilized, with combined datasets representing a growing approach to integrate multiple modalities for enhanced semantic understanding. (3) We summarize the data preprocessing process applied to these various data types, outlining common steps such as data extraction, unqualified data deletion, suitable data representation, and data segmentation into training, validation, and testing sets. (4) To address the challenge of data scarcity in cybersecurity, a notable trend involves leveraging LLMs for data augmentation. Strategies include generating synthetic code samples (vulnerable or benign), specific malicious entities, diverse textual data (like security suggestions or explanations), and assisting in knowledge base construction.
@@ 8.3. Roadmap
Incorporating security knowledge into LLMs. A key direction for the future is to integrate specialized security task solutions and knowledge from the cybersecurity domain into LLMs to overcome potential hallucinations and errors (Aghaei and Al-Shaer, 2023; Ling et al., 2023). This integration aims to enhance LLMs’ ability to address security tasks, especially those requiring a significant amount of domain expertise, such as penetration testing (Deng et al., 2024b; Temara, 2023), hardware vulnerability detection (Dessouky et al., 2019), log analysis (Liu et al., 2023b; Karlsen et al., 2023), and more. Embedding rules and best practices from specific security domains into these models will better represent task requirements, enabling LLMs to generate robust and consistent security task solutions.
@@
Security for Large Language Models (Security4LLM). LLMs have gained considerable traction in the security sector, showcasing their potential in security-related endeavors. Nonetheless, delving into the internal security assessment of LLMs themselves remains a pressing area for investigation (Yao et al., 2024). The intricate nature of LLMs renders them vulnerable to a wide range of attacks, necessitating innovative strategies to fortify the models (Deng et al., 2024a; Liu et al., 2023a; Glukhov et al., 2023; Yao et al., 2024). These vulnerabilities can be broadly categorized (Yao et al., 2024). AI-inherent vulnerabilities stem from the machine learning aspects and include adversarial attacks like data poisoning (manipulating training data to compromise the model (Wan et al., 2023)) and backdoor attacks (embedding hidden triggers), as well as privacy-compromising inference attacks (attribute inference (Staab et al., 2024), membership inference (Mireshghallah et al., 2022)) and extraction attacks (recovering sensitive training data (Carlini et al., 2021)). Instruction tuning attacks, such as jailbreaking (bypassing safety restrictions (Deng et al., 2024a)) and prompt injection (manipulating prompts to elicit unintended behavior (Greshake et al., 2023)), exploit the way LLMs process instructions. Non-AI inherent vulnerabilities may arise from the surrounding infrastructure, such as remote code execution (RCE) in integrated applications (Liu et al., 2024b) or supply chain vulnerabilities involving third-party plugins (Iqbal et al., 2025).
@@ 9. Conclusion
LLMs are making waves in the cybersecurity field, with their ability to tackle complex tasks potentially reshaping many cybersecurity practices and tools. In this comprehensive literature review, we delved into the emerging uses of LLMs in cybersecurity. We firstly explored the diverse array of security tasks where LLMs have been deployed, highlighting their practical impacts (RQ1). Our analysis covered the different LLMs employed in security tasks, discussing their unique traits and applications (RQ2). Additionally, we examined domain-specific techniques for applying LLMs to security tasks (RQ3). Lastly, we scrutinized the data collection and preprocessing procedures, underlining the importance of well-curated datasets in effectively applying LLMs to address security challenges (RQ4). We outlined key challenges facing LLM4Security and provided a roadmap for future research, outlining promising avenues for exploration.

### v5

@@ 3.6. Application of LLM Agents in Security
RQ1 - Summary (1) We have divided cybersecurity tasks into six domains: software and system security, network security, information and content security, hardware security, and blockchain security. We have summarized the specific applications of LLMs in these domains. (2)We discussed 29 cybersecurity tasks and found that LLMs are most widely applied in the field of software and system security, with 119 papers covering 11 tasks. (3) A significant emerging trend identified is the use of LLM-based autonomous agents. Our analysis highlights that these agents represent a paradigm shift, moving from performing isolated security tasks to orchestrating complex, multi-step workflows across all major security domains.
@@ 4.2. Trend Analysis
RQ2 - Summary (1) We have gathered papers utilizing over 30 distinct LLMs for cybersecurity tasks. These LLMs have been categorized into three groups based on their underlying architecture or principles: encoder-only, encoder-decoder, and decoder-only LLMs. (2) We analyzed the trend in employing LLMs for cybersecurity tasks, revealing that decoder-only architectures are the most prevalent. Specifically, over 15 LLMs fall into the decoder-only category, and a total of 133 papers have investigated the utilization of decoder-only LLMs in cybersecurity tasks. (3) Our analysis reveals a clear trend of task-architecture alignment. While decoder-only models have become dominant since 2023 for their generative and reasoning capabilities, encoder-only and encoder-decoder architectures remain highly relevant for specific niches. Encoder-only models are consistently chosen for representation-heavy classification tasks, while encoder-decoder models are often preferred for structured sequence-to-sequence tasks like program repair.
@@ 5.1. Fine-tuning LLMs for Security Tasks
Beyond standard supervised fine-tuning, Reinforcement Learning (RL) based strategies are gaining attention for aligning LLMs more closely with specific security objectives or leveraging their outputs to guide automated processes (Ouyang et al., 2022; DeepSeek-AI et al., 2025; François-Lavet et al., 2018). These approaches often involve optimizing an agent’s policy using reward signals derived from task execution or human feedback, sometimes incorporating LLM-generated insights into the reward function (Tang et al., 2024). For instance, in web application security, RL can fine-tune an LLM to generate more effective web fuzzing inputs by rewarding payload sequences that successfully bypass defenses or trigger desired responses (Liang et al., 2024). Similarly, RL techniques can enhance log analysis by optimizing models based on the accuracy or utility of the identified anomalies (Han et al., 2023). In the domain of fuzzing complex systems like JavaScript engines, CovRL utilizes coverage-guided RL: it processes execution coverage using methods like TF-IDF to assign higher importance to rarely executed paths, then uses this weighted coverage to generate a reward signal, thereby fine-tuning an LLM mutator to produce mutations more likely to explore these less-covered, potentially bug-rich areas (Eom et al., 2024). For automated program repair, RePair leverages RL with process-based feedback; instead of just learning from fixed code examples, the LLM’s policy for generating patches is directly rewarded based on the outcome of executing the generated code against test suites, thus encouraging the model to learn repair strategies that lead to functional correctness (Zhao et al., 2024). In a different application connecting LLMs and RL, CrashTranslator utilizes LLMs first to analyze stack traces and compute semantic similarities with UI elements; this information then shapes the reward function for an RL agent that learns an optimized UI exploration policy to effectively reproduce mobile application crashes (Huang et al., 2024b).
@@ 5.3. External Augmentation
External tools Analysis results from specialized tools serving as auxiliary inputs or validate the result with external tools. Static analysis tools, penetration testing tools, Rule-Based Reasoning Engine (Happe et al., 2023) (Alrashedy and Aljasser, 2024) (Armengol-Estapé et al., 2024) (Cao et al., 2024)(Nazzal et al., 2024) (Xu et al., 2025b)(Chen and Jabbarvand, 2024) (Zhang et al., 2024c)(Zhang et al., 2024d) (Huang et al., 2024b)(Zhang and Zhang, 2024) (Ahmad et al., 2024a)
@@
Inter-model interaction Introducing multiple models (which can be LLMs or other models) to collaborate and interact. Multiple LLMs collaboration, graph neural networks, gGAN (Cai et al., 2024a) (Yang et al., 2024a)(Tang et al., 2023) (Nazzal et al., 2024)(Huang et al., 2024b) (Mündler et al., 2024)(Wang et al., 2024b)
Rebroadcasting Applicable to multi-step tasks, broad-casting the output results of each step iteratively as part of the prompt for the next step. Difficulty-based patch example replay, variables’ name propagation (Xu et al., 2023) (Yuan et al., 2022)(Xu et al., 2025b)
@@
We summarized the external augmentation techniques combined with LLMs in previous studies, as shown in Table 7, identifying 8 distinct categories. The first augmentation technique is feature augmentation. The effectiveness of LLMs in handling downstream tasks heavily relies on the features included in the prompts. We have observed that many studies employing LLMs for security tasks extract contextual relationships or other implicit features from raw data and integrate them with the original data to customize prompts. These implicit features encompass descriptions of vulnerabilities (Zhang et al., 2023f), bug locations (Jin et al., 2023), threat flow graphs (Liu et al., 2023c), and more. Incorporating these implicit features alongside raw data leads to enhanced performance compared to constructing prompts solely from raw data. The next augmentation technique is external retrieval. External knowledge repositories can mitigate the hallucinations or errors arising from the lack of domain expertise in LLMs. LLMs can continually interact with external knowledge repositories during pipeline processing and retrieve knowledge relevant to security tasks to provide superior solutions (Perrina et al., 2023). Next are external tools; rule-based external tools can serve as specialized external knowledge repositories, and in addressing security tasks, LLMs can utilize results from external tools (like static analyzers or formal methods) to rectify their outputs, thereby avoiding redundancy and errors(Happe et al., 2023; Alrashedy and Aljasser, 2024; Xu et al., 2025c). The fourth augmentation technique is task-adaptive training. Existing studies adopt various training strategies (beyond standard fine-tuning discussed in Section 5.1), like contrastive learning or distillation, to strengthen LLMs’ adaptability to complex security tasks, enabling them to generate more targeted outputs (Chen et al., 2020; Tang et al., 2023; Fang et al., 2023). The fifth augmentation technique, inter-model interaction, has garnered significant attention when a single LLM may struggle to handle complex and intricate tasks. Decomposing the pipeline process and introducing multiple LLMs for enhanced performance have been explored (Yang et al., 2024a). This approach leverages collaboration and interaction among models to harness the underlying knowledge base advantages of each LLM. When a single interaction is insufficient to support LLMs in tasks such as variable name recovery or generating complex program patches  (Xu et al., 2023; Yuan et al., 2022), rebroadcasting is used: constructing prompts for LLMs multiple times continuously where the output results of each step are broadcast iteratively as part of the prompt for the next step. This helps reinforce the contextual relationship between each interaction, thereby reducing error rates. The next augmentation technique is post-processing, where LLMs’ outputs are validated or processed for certain security tasks requiring specific types of output (Tihanyi et al., 2023). This process helps mitigate issues such as hallucinations arising from the lack of domain knowledge in LLMs (Chen et al., 2023c). The last strategy is workflow integration, which involves integrating LLMs as components within specific task workflows, incorporating various forms of interactive feedback. This feedback, such as coverage metrics guiding fuzzing strategies (Wang et al., 2024f; Lyu et al., 2024), self-check reasoning steps (Wang et al., 2024f), or outputs being integrated into human rating systems (Thomas et al., 2024), allows the workflow’s context and objectives to steer the LLM’s behavior.
@@
RQ3 - Summary (1) We summarize the domain-specific techniques used in previous research to apply LLMs to security tasks, including prompt engineering, fine-tuning, and external augmentation. (2) Prompt engineering is the most widely used domain technique, with almost all 185 papers employing this approach. Fine-tuning techniques were used in 31% of the papers, while task-specific external augmentation techniques were adopted in 36% of the papers. (3) We categorize and discuss the fine-tuning and external augmentation techniques mentioned in these papers, and analyze their relevance to specific security tasks. (4) A key insight is that state-of-the-art LLM4Security applications are increasingly designed as hybrid systems. While prompt engineering is universal, the most sophisticated approaches specialize the model’s capabilities, either through internal adaptation (fine-tuning) or by integrating it with external data and tools (external augmentation). This indicates a clear trend towards using LLMs as a core reasoning engine within a larger, more reliable security workflow, rather than as standalone oracles.
@@ 6.2. Types of Datasets
Buggy code and bug reports 4 (Zhang et al., 2024c) (Huang et al., 2024b) (Wang et al., 2024c) (Tang et al., 2024)
@@
Decompiled code and variable names 1 (Xu et al., 2025b)
@@
Combined datasets, integrating multiple data types, represent a growing approach (23 instances identified). This strategy often pairs code artifacts with relevant natural language text, such as combining vulnerable code with vulnerability descriptions (Liu et al., 2023d; Chen et al., 2023c), buggy code with bug reports (Zhang et al., 2024c; Huang et al., 2024b), or source code with program documentation (Zhang et al., 2024e). The key advantage lies in leveraging the LLM’s ability to process both structured code and unstructured text simultaneously (Wang et al., 2024a). By providing natural language context (e.g., descriptions of what a vulnerability entails or what a function should do), these combined datasets help LLMs achieve a deeper semantic understanding of the code during fine-tuning or prompting, leading to more accurate vulnerability detection, better code repair suggestions, or more effective code analysis compared to using code data alone.
@@ 6.4. Data Augmentation using LLMs
RQ4 - Summary (1) Our analysis of data sources reveals a significant reliance on open-source, collected, and constructed datasets, reflecting a persistent challenge of data scarcity. A critical insight is the simultaneous disconnect from industrial practice, evidenced by the extreme scarcity of industrial datasets (used in only 1% of papers), which raises questions about the real-world applicability of current models. (2) We categorize datasets into three primary types: code-based, text-based, and combined. While code-based data is predominant, a key trend is the growing use of combined datasets, which pair code with natural language descriptions to provide crucial context and enhance the semantic understanding of LLMs. (3) To address the data scarcity challenge, a notable insight is the trend of leveraging LLMs for data augmentation. Our review shows researchers are increasingly using LLMs to generate synthetic code samples, malicious entities, and diverse textual data to enhance the size and quality of training sets. (4) We summarize the common data preprocessing pipeline applied across studies, which includes essential steps such as data extraction, cleaning and deduplication, converting data into suitable representations for LLMs, and segmentation into training, validation, and testing sets.
@@ 8.3. Roadmap
Incorporating security knowledge into LLMs. A key direction for the future is to integrate specialized security task solutions and knowledge from the cybersecurity domain into LLMs to overcome potential hallucinations and errors (Aghaei and Al-Shaer, 2023; Ling et al., 2023). This integration aims to enhance LLMs’ ability to address security tasks, especially those requiring a significant amount of domain expertise, such as penetration testing (Deng et al., 2024b; Temara, 2023), hardware vulnerability detection (Dessouky et al., 2019), log analysis (Liu et al., 2023b; Karlsen et al., 2023), and more. Embedding rules and best practices from specific security domains into these models will better represent task requirements, enabling LLMs to generate robust and consistent security task solutions (Cheng et al., 2025).
@@
Security for Large Language Models (Security4LLM). LLMs have gained considerable traction in the security sector, showcasing their potential in security-related endeavors. Nonetheless, delving into the internal security assessment of LLMs themselves remains a pressing area for investigation (Yao et al., 2024; Huang et al., 2024a; Das et al., 2025). The intricate nature of LLMs renders them vulnerable to a wide range of attacks, necessitating innovative strategies to fortify the models (Deng et al., 2024a; Liu et al., 2023a; Glukhov et al., 2023; Yao et al., 2024; Zhou et al., 2025). These vulnerabilities can be broadly categorized (Yao et al., 2024). AI-inherent vulnerabilities stem from the machine learning aspects and include adversarial attacks like data poisoning (manipulating training data to compromise the model (Wan et al., 2023)) and backdoor attacks (embedding hidden triggers), as well as privacy-compromising inference attacks (attribute inference (Staab et al., 2024), membership inference (Mireshghallah et al., 2022)) and extraction attacks (recovering sensitive training data (Carlini et al., 2021)). Instruction tuning attacks, such as jailbreaking (bypassing safety restrictions (Deng et al., 2024a)) and prompt injection (manipulating prompts to elicit unintended behavior (Greshake et al., 2023)), exploit the way LLMs process instructions. Non-AI inherent vulnerabilities may arise from the surrounding infrastructure, such as remote code execution (RCE) in integrated applications (Liu et al., 2024b) or supply chain vulnerabilities involving third-party plugins (Iqbal et al., 2025).
@@ 9. Conclusion
The rapid integration of LLMs is creating a paradigm shift in cybersecurity, particularly by revolutionizing how security is integrated into the software engineering lifecycle. In this systematic literature review of 185 seminal papers, we have not only mapped the landscape of LLM4Security but also synthesized several key insights that highlight this transformation:
  • A Paradigm Shift from Analysis to Generation and Repair: A primary contribution of LLMs to software engineering is moving beyond traditional defect detection to automated generation and repair. Decoder-only models, in particular, are being leveraged to create and validate patches for software vulnerabilities and bugs, representing a significant leap towards self-healing software.
  • Solving Data Scarcity through Generative Augmentation: LLMs offer a novel solution to the long-standing problem of data scarcity in security research. Our review found a prominent trend of using LLMs to generate high-quality synthetic data—from vulnerable code snippets to textual repair suggestions—thereby creating richer and more diverse datasets for training robust security tools.
  • The Rise of Hybrid Intelligence Workflows: The most effective applications of LLMs do not operate in isolation. Instead, they act as intelligent ”brains” orchestrating traditional software engineering tools like static analyzers, fuzzers, and verifiers. This hybrid, neuro-symbolic approach, which often involves external augmentation and tool integration, is becoming the de facto standard for building reliable LLM-powered security solutions.
  • The Emergence of Autonomous Agents: Beyond single-task automation, a forward-looking trend is the development of autonomous LLM-based agents capable of handling complex, multi-step security workflows. These agents, which can plan, use tools, and learn from interactions, are being applied to entire vulnerability lifecycles and penetration testing, pointing towards a future of autonomous security operations.
These insights are the culmination of our in-depth analysis guided by four Research Questions (RQs), which systematically examined the security tasks (RQ1), LLM models (RQ2), adaptation techniques (RQ3), and data handling practices (RQ4) in the field. By also outlining the critical challenges and providing a roadmap for future work, this survey serves as a valuable resource for researchers and practitioners aiming to navigate and contribute to the rapidly evolving intersection of large language models and cybersecurity.